logv5 (3x-10)-3logv5 4=2
the "v" means sub and the 4 is after the sub5 not 54
im having a lot of trouble and i could really use someones help :)

Answer by dabanfield(803)

1.) logv5 (3x-10)-3logv5 4 = 2

Using two laws of logs:
log a^b = b*log a and
log (a- b) = log a - log b
We have from 1.):
logv5 (3x-10) - logv5 4^3 = 2 (note that 4^3 = 64)
logv5 (3x-10)/64 = 2
By definition of a logv5 we have:
(3x-10)/64 = 5^2
3x-10 = 25*64
Solve the above for x.